OpenStudy (anonymous):

How many solutions are there to the following system of equations? 2x - y = 2 -x + 5y = 3

OpenStudy (anonymous):

2x - y = 2 2x-2=2 2x=2.2 2x=4 x=4/2 x=2 -2+5y=3 5y=3+2 5y=5 y=5/5=1 x=2 y=1